Sol Summary – December 2nd

Sol: 5
Summary Title: Taking Flight
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: EVA excursions in the local area for the first drone flight, to retrieve the mirror experiment for interim examination in the Science Dome, and, finally, to re-deploy the mirror experiment.
Look Ahead Plan: EVAs to conduct soil samples, working from Somerville Overlook and Cowboy Corner.
Anomalies in work: None
Weather: High winds in the morning, with warmer than typical temperatures. Winds slowly eased throughout the day, making for a calm sunset under clear skies.
Crew Physical Status: Crew is in good shape, nothing of concern to note.
EVA: EVA #1 led by the XO and Crew Engineer on foot to retrieve the mirror experiment, 1 hour duration. EVA #2 led by the Technology Officer, with support from the Commanding Officer and Health and Safety Officer. 2 1/2 hour duration using Spirit and Perseverance rovers.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

Commander Report – December 2nd

Sol: 5
Summary Title: EVAs #5 and #6
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. Winds were moderate most of the day, with higher gusts, with a brief drizzle in the late morning. The crew enjoyed a breakfast of cereal, oatmeal, homemade sourdough bread, and espresso. After a morning planning meeting, HSO collected crew health stats. XO and Engineer conducted a short morning EVA (#5) to retrieve the mirror samples to analyze in the Science Dome under a microscope. Technology Officer employed help from the two licensed private/drone pilots on the crew (CDR, HSO) to complete drone flight planning preparations for the afternoon EVA. After a light lunch, CDR, Technology Officer, and HSO embarked on EVA #6, while the Green Hab Officer led IVA Comms. The EVA plan included drone operations with flight plans near the Hab, near the mirror sample deployment location, and at Marble Ritual. The crew successfully completed activation and checkout of the drone near the Hab but decided to call off Flight Plan #1 due to high wind gusts. The crew proceeded to perform their backup EVA plan operations demonstrating the VR/AR headset in the area around the Hab and at the intersection of Hab Rd and Cow Dung Rd. At the mirror deployment location and Marble Ritual, the crew again called off the drone Flight Plans #2 and #3 due to high wind gusts and performed further demonstration of the VR/AR headset in those areas. The crew then returned early to the Hab area; by that time the winds had died down slightly allowing for the crew to proceed with Flight Plan #1 around the Hab after requesting permission from Tier 1 support to modify the original EVA plan. The drone flight was successful, collecting imagery of the Hab area. Finally, the EVA crew retrieved the mirror samples from the airlock and redeployed them at the location across Cow Dung Rd. from Marble Ritual before returning to the Hab. During the EVA, Engineer analyzed the mirror samples under the Aerospace microscope in the Science Dome. In the afternoon, Green Hab Officer harvested carrot sprouts and the crew continued their montage painting. The crew ended the day with a pizza party.

Operations Report – December 2nd

SOL: 5

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 201.9 (Start), 202.0 (End)
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: Yes
Hours: 250.3 (Start), 250.5 (End)
Beginning charge: 100%
Ending charge: 87%
Currently charging: No

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 45 gallons used today, 155 gallons used in total thus far
Water (static tank): 395 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The crew used the two-piece EVA suits on the EVAs today. Before the first EVA, our Executive Officer, Alli Taylor, first put on Suit #6 and everything was working fine, but before she left the habitat, she noticed that she was getting warm and her helmet started to fog up. It seems that the fan had a very low air flow despite turning up the air flow dial, so she switched suits and took Suit #4 instead.

During the second EVA, our Health and Safety Officer, Barbara Braun, used Suit #4, as well. By the end of the EVA, she was experiencing low air flow. She asked another crew member to turn up her air flow dial, but by the time she returned to the front air lock at the end of the EVA, there was no air flow at all.

From a Crew Engineer’s perspective, I suspect this may be a charging issue because upon inspection of the suits tonight, I noticed that the suits in question were not set to the “off” mode and hence were not charging. I believe this is what caused the lack of airflow, but I will monitor the situation and report back tomorrow.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Today, the GreenHab Officer harvested some carrot sprouts for dinner. Additional information can be found in the GreenHab Report.
WATER USE: ~28 gallons (~5 gallons today)
Heater: On
Supplemental light: Off
Harvest: 28.3 grams (28.3 oz; converted using 1 oz = 28.4 g)

Summary of ScienceDome operations: The Crew Engineer performed mirror inspections using the Aerospace Corp microscope and took images of the abnormalities/anomalies observed.
Dual split: No heat or AC used. All systems off.

Summary of RAM operations: N/A

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The Crew Engineer has not yet installed the furnace air filter. Please advise if this is something you would like us to do tonight/tomorrow. Crew on second EVA today did an external inspection of the habitat. Everything looked good.

Journalist Report – December 2nd

Sol: 5
Author: Matthew Eby, GreenHab Officer

Gale force winds arrived overnight, buffeting the habitat. At 3am the mind wanders: what was the fail-safe wind level? Optimizing for mass, a safety factor, this far, but no further. The Hab swayed and creaked with the gusts until dawn arrived and the wind started to ease just a bit. Today a pair of coordinated EVAs deployed teams across the station, equipment and test samples passing back and forth through the airlock as the science dome and the EVA crews worked in tandem. The XO, Alli Taylor, and the Crew Engineer, Ashley Kowalski, kicked things off with a morning trek across the East flat to retrieve the mirror coating exposure experiment package, returning it for an interim inspection under the microscope at the Science Dome. While this inspection progressed, a second EVA was dispatched with the Commander, Kristine Ferrone, the Health and Safety Officer, Barbara Braun, and the Technology Officer, Trevor Jahn, to attempt the first drone flight. Wind levels threatened to scrub the flight, but the wind relaxed just enough at the end of the EVA window to allow the drone to take flight. Ball’s Bluff Elementary chimed in on our time-delayed STEM outreach channel, christening the drone “Vixl”. Soaring 100 feet above the station, Vixl zig-zagged above the Hab capturing data destined for the augmented reality project. Overcoming hurdles in the dome, the mirror inspection was completed in the nick of time, and the experiment package was returned to the outside environment for additional dust exposure.

The entire crew reunited by the late afternoon. The wind and the crew grew tired and quiet descended across the station. It was time for the first naps of the rotation, quiet time in the GreenHab thinning carrot seedlings, and recording our visions on the group canvas. Recovered, we gathered for a pizza party.

GreenHab Report – December 2nd

GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 77-deg F at time of watering. Range prior 24 hours: 65-deg F – 82-deg F
Hours of supplemental light: 0 hours
Daily water usage for crops: Approximately 5 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 220 gallons (Based on a 250 gallon capacity, visually 90% full)
Time(s) of watering for crops: 4 PM
Changes to crops: None
Narrative: Today some thinning of the existing carrot bed was carried out, it will require additional thinning in a few days. Also of note, the bean plants are close to flowering.
Harvest: 28.3 grams of carrot sprouts harvested as part of thinning effort.
Support/supplies needed: None needed at this time.

EVA Report – December 2nd

EVA # 5
Author: Trevor Jahn, Technology Officer
Purpose of EVA: Retrieve Mirror Samples
Start time: 9:10 am
End time: 9:55 am
Narrative: Allison Taylor and Ashley Kowalski left the airlock and continued on foot to the Mirror Sample location, retrieved the Mirror Samples, and returned to the airlock.
Coordinates (use UTM WGS 84): 518500, 4251000
Participants: Allison Taylor and Ashley Kowalski
Road(s) and routes per MDRS Map: N/A, path of least resistance to and from Mirror Samples, North of the Crew Habitat
Mode of travel: Walking

EVA # 6
Author: Trevor Jahn, Technology Officer
Purpose of EVA: Demonstrate Augmented Reality Programs, and Drone Flight Plan 1, 2 and time permitting Flight Plan 3.
Start time: 12:00 pm
End time: 2:30 pm
Narrative: Trevor, Kristine, and Barbra left the airlock at 12:00 pm for the second EVA of the day and proceeded on foot clockwise around the crew habitat until they were approximately 100 ft due north of the crew habitat. The three began their drone Flight Plan 1 launch procedures, but before launching the drone to proceed along it’s predetermined path, the crew called off the flight due to wind conditions being too variable for a maiden flight and proceeded on foot back to the front of the crew habitat, on the crew habitat’s south side, where they deposited their equipment and conducted an Augmented Reality (AR) software demonstration. The trio proceeded, after their AR demonstration, in the crew rovers, Perseverance and Spirit, to the corner of Cow Dung Road and Habitat Road, conducted a second AR demonstration, before continuing onto Marble Ritual, where they left the crew rovers, and continued foot to the Mirror Sample deposit location. The Mirror Samples were not there by design but had been collected during an EVA earlier in the morning, but the crew conducted another AR demonstration at the Mirror Sample location, before assessing the weather conditions to determine if they should proceed with drone Flight Plan 2 and 3 but concluded the weather conditions had not improved enough to continue with the flight. The three astronauts continued to investigate the flight area of drone Flight Plan 3, which was over Marble Ritual, followed by continuing back to the crew rovers, and returning to the crew habitat. The Mirror Samples, which were retrieved during the morning EVA, were not ready to be returned to the airlock for the EVA crew to return to their deployment location, so a request was put in to Tear 1 Mission Support, to make a second attempt at drone Flight Plan 1. This request was granted, and the EVA crew conducted a successful EVA flight. Mirror Samples were deployed in the airlock, and the EVA crew returned them to their deployment location on foot and returned to the Crew Habitat airlock without incident, ending their EVA at 2:30 pm, 30 min before their 3:00 pm deadline.
Destination: Marble Ritual and Galileo Road
Coordinates (use UTM WGS 84): 518100, 4251000, 519000, 4252000 and 520000, 4251500
Participants: Dr. Kristine Ferrone, Barbara Braun, Trevor Jahn
Road(s) and routes per MDRS Map: Cow Dung Rd (0110) to Marble Ritual
Mode of travel: Rovers (Perseverance and Spirit), Walking

Supplemental Operations Report 03Dec2022

[title Supplemental Operations Report – December 3rd]

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Generator

Action taken for non-nominal systems: Propane regulator was checked and pressure increased from 9 to 11. The generator seems to be running better, but there is still some sort of miscommunication between the generator and power system. Still working on it.

Generator: Running from about 8 pm to 8 am.

ScienceDome Dual Split: Heat on between 8 pm and 8 am at 67degrees F.

Solar— Nominal.
Solar— SOC Last 24 hours:
Max 100%
Min 52%
Avg 70.0% (under heavy cloud cover)
Propane Reading, station tank – filled today %
Propane Reading, director tank— filled today %
Propane Reading, intern tank— filled today %
Propane Reading, generator— filled today %

Ethanol Free Gasoline – 0 gallons

Water (static tank) – 175 gallons

Water in GreenHab – Not noted gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 175 gallons

Hab toilet tank emptied: check regular operations reports

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We disconnected the heat sensor from the motor in Oppy. Oppy has had a faulty sensor since it was new (this is a factory problem for some Polaris Ranger EVs). As a result, Oppy has been underused. The only way to fix the problem was to replace the motor, which would cost $2800. Instead we disconnected the sensor, then tested the rover by driving it as fast and as safely as possible to the cattle guard at the beginning of Cow Dung Road and back. We checked the motor temperature by touch several times and each time it was only as hot as Soji. The only downside to this repair is that if the motor does actually overheat in the future, the motor will fail. Then we will have to replace it, but if this repair holds we will have a fully functioning rover. It currently has less than half the hours of every other rover, and now it can be used more.

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda

Reason for use: charging battery

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Currently waiting to be picked up for serving. After I started and moved the Honda today, it got stuck in 2nd gear. This means that none of the ATVs are currently operational.

HabCar used and why, where? Yes, to town and to the dump

CrewCar used and why, where? Yes, to town and to Grand Junction

Luna used and why, where? Yes, to town and to Loa

General notes and comments: Tabs for the year were purchased for the 2018 rovers and the HabCar. That means all vehicles are registered until next year.

Summary of internet: Nominal. HudgesNet is currently off, and we plan to reroute it for crew use and keep Starlink at the Outpost except when crews are out of sim. The new iMac has been installed at the HabCom Station on the lower deck, however we will be moving it back to the Outpost and relocating the desk back to the SciDome when this crew is completed. We need to reimagine the HabCom station because where we put it, it was just getting used as a workbench.

Summary of suits and radios: See crew operations reports. The next upgrade to the repeater system will take place when Crew 271 is on station.

Campus wide inspection, if action taken, what and why? There are great advantages to having two people working on campus. Today we hitched the trash trailer up to the HabCar and delivered years of accumulated waste to the dump. The trailer still has the one tire with the slow leak but otherwise worked well. We added air to all tires so that each was at 30 psi.

Summary of general operations: Started work on moving the conduit.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Plants were fertilized. Tomato cages were installed. Timer was installed for four hours of supplemental light from 10 pm – 2 am.

Summary of ScienceDome operations: Nothing to report. Duel split turned on to 67 degrees F overnight, same timing as the generator.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: n/a

Copyright © The Mars Society. All rights reserved. | Main Site